# Positive Text Reframing under Multi-strategy Optimization

### v1

@@ 1 Introduction
The concept of style transfer initially emerges within the domain of computer vision (CV) with the objective of accomplishing image style transfer (Gatys et al., 2016). Inspired by this, Hu et al. (2017) proposed text style transfer (TST), whose main purpose is to automatically control the text style and preserve the style-independent content. There also have been some related research before this, such as paraphrase (Xu et al., 2012). In recent years, there has been an increasing focus on TST, which has gradually evolved into a significant subfield within the domain of natural language generation. Many corresponding task variants also have been proposed, such as text form transfer (Briakou et al., 2021), topic transfer (Huang et al., 2020), text simplification (Cao et al., 2020), and sentiment transfer (Mueller et al., 2017), etc.
@@
$\bullet$ Extensive experimental results demonstrate that our proposed multi-strategy optimization framework achieves significant improvement on both unconstrained and controlled positive reframing task.
@@ 2 Related Work
Early research on text style transfer mostly relied on artificial design features such as syntax (Zhu et al., 2010) and phrase (Xu et al., 2012) modeling, etc. Similar to other tasks in NLP, the advent of deep learning has resulted in the growing application of neural network models to TST. For example, Jhamtani et al. (2017) investigated the utilization of the Seq2Seq model for transforming modern English into Shakespearean-style English. Wang et al. (2019) applied GPT-2 to accomplish the formal-informal transfer. Sancheti et al. (2020) extended the work of Jhamtani et al. (2017) by incorporating a reinforcement learning framework. Lai et al. (2021) further applied this framework to PLMs. Above studies are mainly based on parallel corpora. Although satisfactory results can be achieved, the cost of constructing parallel corpora is expensive. Therefore, semi-supervised learning and unsupervised learning are widely used in TST. The main methods include data augmentation or text retrieval (Zhang et al., 2020; Jin et al., 2019), adversarial learning (Hu et al., 2017; Fu et al., 2018), back-translation (Prabhumoye et al., 2018; Wei et al., 2023), and reinforcement learning (Luo et al., 2019; Gong et al., 2019).
@@ 3.1 Problem Definition
Regarding reframe strategy classification, its requirement is to predict the positive reframing strategy $\psi_{x}$ used to reframe the original sentence $x$.
@@ 4.1 Dataset
Label Train Dev Test
Growth 1683 216 221
Impermanence 1296 172 157
Neutralizing 2410 303 302
Optimism 3295 373 400
Self-affirmation 673 92 76
Thankfulness 882 94 109
@@ 4.2 Evaluating Metrics
For generation task, the following nine automatic metrics are used: (1) Content preservation-related metric, namely ROUGE-1 (R-1), ROUGE-2 (R-2), ROUGE-L (R-L) (Lin, 2004), BLEU (Papineni et al., 2002) and BERTScore (BScore) (Zhang et al., 2019). (2) $\Delta$TextBlob ($\Delta$TB) (Loria, 2018) is used to report the average change in sentiment. (3) RTQE (Reframing Text Quality Evaluation) is proposed to evaluate the degree of positive text reframing (i.e. style strength), we fine-tune RoBERTalarge (Liu et al., 2019) to evaluate reframing degree and we regard the probability from the model prediction as the degree of positive reframing between the original and generated sentence; on the human reference it has the F1 score of 95.98% and accuracy of 97.41%. (5) Perplexity (PPL) is an indicator of text fluency, and we use GPT-2large as the evaluation model.
@@ 4.3 Implementation Details
Positive reframing Following Ziems et al. (2022), we use T5 (Raffel et al., 2019) and BART (Lewis et al., 2020) with 6 layers in each of the encoder and decoder, and the hidden size of 768. The value of the learning rate is from 3e-5 to 3e-4, the batch size processed by each device is 6, and the text maximum input length is 80. $\alpha$, $\beta$, $\gamma$ are respectively set to 1, 0.2, 1.
@@ 4.4.1 Reframe Strategy Classification
In addition, the performance of the input approach of directly connecting the original and generated sentence is also tested to demonstrate the effectiveness of the contextual semantic enhancement strategy (i.e., the construction of auxiliary question) used in this paper. And the experimental results are given in Table 4. As can be seen, the F1 score on each label is greatly reduced without context enhancement strategy, but our models still achieve comparable performance with the multi-label classification models which once again proves the effectiveness of our method.
@@
BART (Ziems et al., 2022) 27.7 10.8 24.3 10.3 89.3 0.23 63.8 86.0
@@
Table 5: The experimental results of unconstrained positive reframing. The best in-category performance is bolded and the best overall performance is highlighted. And expect for PPL, all other metrics are better when they are higher.
@@ 4.4.2 Unconstrained Positive Reframing
As shown in Table 5, our proposed framework MSOF achieves significant improvements compared to the baselines. When combining positive sentiment reward and content preservation reward only during the training process, i.e. MSOFGreedy, already outperforms the baselines on almost all metrics, especially ROUGE, BScore, RTQE, and PPL. When incorporating decoding optimization and multi-dimensional re-ranking, the performance of the model will be further improved. From the perspective of the model, the T5-based models achieve the best results on metrics such as $\Delta$TB, RTQE and PPL, while the BART-based models reach SOTA on content preservation-related metrics such as ROUGE, BLEU, and BScore. This may be because BART prioritizes semantic preservation rather than sentiment change when reframing the negative text. Among different decoding methods, both beam search and random sampling-based methods are superior to greedy search. Specifically, Top-k sampling has the best overall performance, achieving the best or sub-optimal results on almost all metrics. Top-p sampling performs slightly lower than Top-k sampling. Compared to the above two decoding methods, beam search and Typical sampling are not satisfactory but still superior to the baseline method. Ultimately, regardless of whether T5 or BART is used as the basic generation model, MSOFTop-k achieves the best results among all variant models, basically achieving at least 7% improvement on each metric compared to baselines, which strongly proves the effectiveness of our proposed framework.
@@
BART (Ziems et al., 2022) 28.8 10.9 25.1 10.1 89.6 0.27 69.5 89.1
@@ 4.4.3 Controlled Positive Reframing
Since only Ziems et al. (2022) have studied controlled positive reframing, we use T5 and BART Ziems et al. (2022) that are fine-tuned on the corresponding dataset as baselines for comparison. The primary experimental results are given in Table 6. It can be concluded that the performance of models under constraints is generally better than unconstrained, which proves that the reframe strategy plays a role in assisting model inference to a certain extent. Consistent with the experimental results under the unconstrained setting, MSOFTop-k still achieves the best results among all variant models. Compared with the baselines, MSOFTop-k achieves an average improvement of 5 points on ROUGE, 1 point in BLEU, more than 10 points on both RTQE and PPL, and an improvement of about 20% on $\Delta$TB. Moreover, it can be found that although Typical sampling does not perform as well as other decoding approaches on content preservation-related metrics such as ROUGE, BLEU, and BScore, it still achieves impressive results on $\Delta$TB, RTQE and PPL, suggesting that its corresponding output is consistent with task requirements to some extend, even though there is less overlap with human reference.
@@ 4.4.4 Ablation Experiment
In addition, from the ablation experimental results shown in Table 7, we can conclude that only applying content preservation reward helps the model perform well on ROUGE, BLEU and BScore, but hinders the model from transferring text style. When using only positive sentiment reward, although the model performs well on $\Delta$TB and RTQE, it is not satisfactory in terms of content preservation. However, when the two are combined, the model can achieve a better balance between sentiment change and content preservation, exhibiting a more comprehensive performance. Furthermore, it can be observed that the multi-dimensional re-ranking significantly improves the model’s performance on multiple metrics. This demonstrates that it can effectively select the sentence from the candidate that better meets the requirements of positive reframing. Based on the above experimental results and analysis, the validity and rationality of each component of MSOF can be effectively proved. For more ablation experiments, please refer to Tables 13 and 14 in Appendix D.2 and Tables 15, 16 and 17 in Appendix D.3.
@@ A.5 Experiment Results
It can be seen from Table 10 that the performance of RoBERTa is generally better than BERT on all metrics, and the large version is better than the base, which proves that the more parameters and training corpus the model has, the better its performance will be. In the end, RoBERTalarge basically achieves the best results in all metrics and also reaches the F1 score of 95.98% and accuracy of 97.41% in the test of evaluating human reference, so finally this paper uses it as the evaluation model for RTQE.
@@ D.3 Controlled Positive Reframing
Additionally, we present the ablation results of multi-dimensional re-ranking under controlled setting in Table 17. It can be observed that when the strategy consistency evaluation is not used, the scores of MSOFTop-k on RTQE and PPL will decrease significantly, but it has better performance on ROUGE and BLEU. When the text similarity evaluation is not used, the performance of MSOFTop-k would significantly lower on content preservation-related metrics, but achieves best or sub-optimal results on $\Delta$TB and RTQE. And when the fluency evaluation is not used, the model scores significantly lower on PPL, but still achieves sub-optimal results on RTQE and content preservation-related metrics. This paper suggests that the reason for the above phenomenon may be that the strategy consistency evaluation considers excessive content preservation as indicating incomplete reframing, and thus interacts with the text similarity evaluation. In addition, as can be seen from the results in the table, a decrease in text fluency (high PPL) is often accompanied by a decrease on $\Delta$TB and RTQE. Therefore, there may be some positive correlation among them. Finally, although the overall framework does not achieve optimal results on all metrics, considering the performance of each variant model on each metric, choosing this way is the best trade-off at present.
D.4 Case Study
@@
We provide the generated examples of unconstrained and controlled experiments in Tables 18 and 19. A comparative analysis reveals that our models generate outputs that are more diverse and comprehensive, while effectively preserving the underlying meaning of the original text. Specifically, the outputs of the BART-based models are mostly similar, except for the sentences generated by Typical sampling. On the other hand, the T5-based models outperform the BART-based models and baselines by providing the benefits of weekends consistent with human reference. Additionally, although the text in the dataset may contain colloquialisms and even grammatical errors, our models can generate more formal sentences that avoid these issues. Therefore, we speculate that further cleaning and filtering of the data in the dataset can further improve the model’s performance. By comparing the results generated by the model in the unconstrained and controlled settings, it can be inferred that without reframe strategy, the reframing performance of the models will decrease, which proves that the reframing strategy plays an auxiliary role in helping the model generate results that better meet task requirements.
@@
Table 20: A model comparison for reframing the same text using different reframe strategy

### v3

@@ Abstract
Positive Text Reframing under Multi-strategy Optimization
Shutong Jia1,2, Biwei Cao1, Qingqing Gao1, Jiuxin Cao1thanks:   Corresponding author., Bo Liu1 1Southeast University, 2State Grid Tianjin Power Dongli Power Supply Branch {shutong_jia,caobiwei,qingqing_gao,jx.cao,bliu}@seu.edu.cn
@@ 1 Introduction
The concept of style transfer initially emerges within the domain of computer vision (CV) with the objective of accomplishing image style transfer (Gatys et al., 2016). Inspired by this, Hu et al. (2017) proposed text style transfer (TST), whose main purpose is to automatically control the text style and preserve the style-independent content. In recent years, there has been an increasing focus on TST, which has gradually evolved into a significant subfield within the domain of natural language generation. Many corresponding task variants also have been proposed, such as text form transfer (Briakou et al., 2021), topic transfer (Huang et al., 2020), text simplification (Cao et al., 2020), and sentiment transfer (Mueller et al., 2017), etc.
@@
$\bullet$ Extensive experimental results demonstrate that our proposed multi-strategy optimization framework achieves significant improvement on both unconstrained and controlled positive reframing task. And we would release our code to encourage future research [^1] https://github.com/20174376/code-for-paper.
@@ 2 Related Work
Early research on text style transfer mostly relied on artificial design features such as syntax (Carroll et al., 1999) and phrase (Quirk et al., 2004) modeling, etc. Similar to other tasks in NLP, the advent of deep learning has resulted in the growing application of neural network models to TST. For example, Jhamtani et al. (2017) investigated the utilization of the Seq2Seq model for transforming modern English into Shakespearean-style English. Wang et al. (2019) applied GPT-2 to accomplish the formal-informal transfer. Sancheti et al. (2020) extended the work of Jhamtani et al. (2017) by incorporating a reinforcement learning framework. Lai et al. (2021) further applied this framework to PLMs. Above studies are mainly based on parallel corpora. Although satisfactory results can be achieved, the cost of constructing parallel corpora is expensive. Therefore, semi-supervised learning and unsupervised learning are widely used in TST. The main methods include data augmentation or text retrieval (Zhang et al., 2020; Jin et al., 2019), adversarial learning (Hu et al., 2017; Fu et al., 2018), back-translation (Prabhumoye et al., 2018; Wei et al., 2023), and reinforcement learning (Luo et al., 2019; Gong et al., 2019).
@@ 3.1 Problem Definition
Regarding reframe strategy classification, its requirement is to predict the reframing strategy $\psi_{x}$ used to reframe the original sentence $x$.
@@ 4.1 Dataset
Label Train Dev Test
Growth 1683 216 221
Impermanence 1296 172 157
Neutralizing 2410 303 302
Optimism 3295 373 400
Self-affirmation 673 92 76
Thankfulness 882 94 109
@@ 4.2 Evaluating Metrics
For generation task, the following nine automatic metrics are used: (1) Content preservation-related metric, namely ROUGE-1 (R-1), ROUGE-2 (R-2), ROUGE-L (R-L) (Lin, 2004), BLEU (Papineni et al., 2002) and BERTScore (BScore) (Zhang et al., 2019). (2) $\Delta$TextBlob ($\Delta$TB) (Loria, 2018) is used to report the average change in sentiment. (3) RTQE (Reframing Text Quality Evaluation) is proposed to evaluate the degree of positive text reframing (i.e. style strength), we fine-tune RoBERTalarge (Liu et al., 2019) to evaluate reframing degree and we regard the probability from the model prediction as the degree of positive reframing between the original and generated sentence; on the human reference it has the F1 score of 95.98% and accuracy of 97.41%. For more details refer to Appendix A (4) Perplexity (PPL) is an indicator of text fluency, and we use GPT-2large as the evaluation model.
@@ 4.3 Implementation Details
Positive reframing Following Ziems et al. (2022), we use T5 (Raffel et al., 2019) and BART (Lewis et al., 2020) with 6 layers in each of the encoder and decoder, and the hidden size of 768. The value of the learning rate is from 3e-5 to 3e-4, the batch size processed by each device is 6, and the text maximum input length is 80. $\alpha$, $\beta$, $\gamma$ are respectively set to 1, 0.2, 1, which are the choices obtained from multiple experiments. And the approach of obtaining the candidate sentence set can be seen in Appendix B.
@@ 4.4.1 Reframe Strategy Classification
In addition, the performance of the input approach of directly connecting the original and generated sentence is also tested to demonstrate the effectiveness of the contextual semantic enhancement strategy (i.e., the construction of auxiliary question) used in this paper. And the experimental results are given in Table 4. As can be seen, the F1 score on each label is greatly reduced without context enhancement strategy, but our models still achieve comparable performance with the multi-label classification models which once again shows the effectiveness of our method.
@@
BART Vallina Fine-tune (Ziems et al., 2022) 27.7 10.8 24.3 10.3 89.3 0.23 63.8 86.0
@@
Table 5: The experimental results of unconstrained positive reframing. The best in-category performance is bolded and the best overall performance is highlighted. And except for PPL, all other metrics are better when they are higher.
@@ 4.4.2 Unconstrained Positive Reframing
As shown in Table 5, our proposed framework MSOF achieves significant improvements compared to the baselines. When combining positive sentiment reward and content preservation reward only during the training process, i.e. MSOFGreedy, already outperforms the baselines on almost all metrics, especially ROUGE, BScore, RTQE, and PPL. When incorporating decoding optimization and multi-dimensional re-ranking, the performance of the model will be further improved. From the perspective of the model, the T5-based models achieve the best results on metrics such as $\Delta$TB, RTQE and PPL, while the BART-based models reach SOTA on content preservation-related metrics such as ROUGE, BLEU, and BScore. This may be because BART prioritizes semantic preservation rather than sentiment change when reframing the negative text. Among different decoding methods, both beam search and random sampling-based methods are superior to greedy search. Specifically, Top-k sampling has the best overall performance, achieving the best or sub-optimal results on almost all metrics. Top-p sampling performs slightly lower than Top-k sampling. Compared to the above two decoding methods, beam search and Typical sampling are not satisfactory but still superior to the baseline method. Ultimately, regardless of whether T5 or BART is used as the basic generation model, MSOFTop-k achieves the best results among all variant models, basically achieving at least 7% improvement on each metric compared to baselines, which strongly indicates the effectiveness of our proposed framework.
@@
BART Vallina Fine-tune (Ziems et al., 2022) 28.8 10.9 25.1 10.1 89.6 0.27 69.5 89.1
@@ 4.4.3 Controlled Positive Reframing
Since only Ziems et al. (2022) have studied controlled positive reframing, we use T5 and BART Ziems et al. (2022) that are fine-tuned on the corresponding dataset as baselines for comparison. The primary experimental results are given in Table 6. It can be concluded that the performance of models under constraints is generally better than unconstrained, which provides support for the reframe strategy to play a role in assisting model inference to a certain extent. Consistent with the experimental results under the unconstrained setting, MSOFTop-k still achieves the best results among all variant models. Compared with the baselines, MSOFTop-k achieves an average improvement of 5 points on ROUGE, 1 point in BLEU, more than 10 points on both RTQE and PPL, and an improvement of about 20% on $\Delta$TB. Moreover, it can be found that although Typical sampling does not perform as well as other decoding approaches on content preservation-related metrics such as ROUGE, BLEU, and BScore, it still achieves impressive results on $\Delta$TB, RTQE and PPL, suggesting that its corresponding output is consistent with task requirements to some extend, even though there is less overlap with human reference.
@@ 4.4.4 Ablation Experiment
In addition, from the ablation experimental results shown in Table 7, we can conclude that applying content preservation reward helps the model perform well on ROUGE, BLEU and BScore, but hinders the model from transferring text style. When using positive sentiment reward, although the model performs well on $\Delta$TB and RTQE, it is not satisfactory in terms of content preservation. However, when the two are combined, the model can achieve a better balance between sentiment change and content preservation, exhibiting a more comprehensive performance. Furthermore, it can be observed that the multi-dimensional re-ranking significantly improves the model’s performance on multiple metrics. This demonstrates that it can effectively select the sentence from the candidate that better meets the requirements of positive reframing. Based on the above experimental results and analysis, the validity and rationality of each component of MSOF can be effectively illustrated. For more ablation experiments, please refer to Tables 13 and 14 in Appendix D.2 and Tables 15, 16 and 17 in Appendix D.3.
@@ 5 Conclusion
Acknowledgement
This work is supported by National Natural Science Foundation of China under Grants No.62172089, No.62472092, No.62106045. Natural Science Foundation of Jiangsu Province, China under Grants No.BK20241751. Jiangsu Provincial Key Laboratory of Computer Networking Technology, China. Jiangsu Provincial Key Laboratory of Network and Information Security, China under Grants No.BM2003201, and Key Laboratory of Computer Network and Information Integration of Ministry of Education of China under Grants No.93K-9, Nanjing Purple Mountain Laboratories, China. Start-up Research Fund of Southeast University under Grants No.RF1028623097. We thank the Big Data Computing Center of Southeast University for providing the facility support on the numerical calculations.
@@ A.5 Experiment Results
It can be seen from Table 10 that the performance of RoBERTa is generally better than BERT on all metrics, and the large version is better than the base, which shows that the more parameters and training corpus the model has, the better its performance will be. In the end, RoBERTalarge basically achieves the best results in all metrics and also reaches the F1 score of 95.98% and accuracy of 97.41% in the test of evaluating human reference, so finally this paper uses it as the evaluation model for RTQE.
@@ D.3 Controlled Positive Reframing
Additionally, we present the ablation results of multi-dimensional re-ranking under controlled setting in Table 17. It can be observed that when the strategy consistency evaluation is not used, the scores of MSOFTop-k on RTQE and PPL will decrease significantly, but it has better performance on ROUGE and BLEU. When the text similarity evaluation is not used, the performance of MSOFTop-k would significantly lower on content preservation-related metrics, but achieves best or sub-optimal results on $\Delta$TB and RTQE. And when the fluency evaluation is not used, the model scores significantly lower on PPL, but still achieves sub-optimal results on RTQE and content preservation-related metrics. This paper suggests that the reason for the above phenomenon may be that the strategy consistency evaluation considers excessive content preservation as indicating incomplete reframing, and thus interacts with the text similarity evaluation. In addition, as can be seen from the results in the table, a decrease in text fluency (high PPL) is often accompanied by a decrease on $\Delta$TB and RTQE. Therefore, there may be some positive correlation among them. Finally, although the overall framework does not achieve optimal results on all metrics, considering the performance of each variant model on each metric, choosing this way is the best trade-off.
D.4 Positive Reframing under ICL
In this section, we test the experimental results of GPT-3.5 on the unconstrained positive reframing task and compare it with MSOF [^3] We use GPT-3.5-turbo-0613 for the experiment. And due to the difficulty of understanding the true meaning of the reframing strategy in LLM under In-Context Learning (ICL), we only conduct experiments under unconstrained setting.. Under zero-shot, we use Rephrase the above sentence to be more positive (Ziems et al., 2024) as the instruction. And under few-shot, following (Sharma et al., 2023; Ziems et al., 2022), we retrieve 5 representative exemplars with the closest semantic similarity from the training set as the context for each test case.
As shown in Table 18, MSOF still outperforms LLM regarding ROUGE, BLEU, BScore, PPL, and $\Delta$TB, but LLM achieves higher RTQE scores. Based on the examples in Table 19, it can be seen that although the output of LLM reflects the concept of reframing, it tends to generate longer text and is more prone to hallucinations. This demonstrates the applicability of MSOF for positive reframing and the continued significance of studying small models in this task.
D.5 Case Study
@@
We provide the generated examples of unconstrained and controlled experiments in Tables 19 and 20. A comparative analysis reveals that our models generate more diverse and comprehensive outputs while effectively preserving the underlying meaning of the original text. Specifically, the outputs of the BART-based models are mostly similar, except for the sentences generated by Typical sampling. On the other hand, the T5-based models outperform the BART-based models and baselines by providing the benefits of weekends consistent with human reference. Additionally, although the text in the dataset may contain colloquialisms and even grammatical errors, our models can generate more formal sentences that avoid these issues. Therefore, we speculate that further cleaning and filtering of the data in the dataset can further improve the model’s performance. By comparing the results generated by the model in the unconstrained and controlled settings, it can be inferred that without reframe strategy, the reframing performance of the models will decrease, which proves that the reframing strategy plays an auxiliary role in helping the model generate results that better meet task requirements.
@@
Setting R-1 R-2 R-L BLEU BScore $\Delta$TB RTQE PPL
Zero-shot 27.0 7.2 21.2 4.2 88.9 0.43 99.8 42.0
Five-shot 28.5 8.5 22.5 5.3 89.3 0.28 97.1 27.5
Table 18: The performance of GPT3.5 on positive reframing.
@@
GPT-3.5 (zero-shot) I’m excited that tomorrow is finally Friday! This week has been full of new experiences and opportunities.
GPT-3.5 (five-shot) I can’t wait for Friday to finally arrive! This week has been a challenge, but I made it through. tired
@@
Table 21: A model comparison for reframing the same text using different reframe strategy.